OpenStudy (anonymous):

solve equation for m 6m+3/7=5m-6/3

OpenStudy (saifoo.khan):

\[6m -5m =- \frac37 -\frac63\]

OpenStudy (saifoo.khan):

\[m = -\frac{17}{7}\]

OpenStudy (kira_yamato):

Take a look at the pic. Remember to remove the denominator before solving

OpenStudy (anonymous):

thanks heaps, that was a great help

OpenStudy (anonymous):

(6m+3)/7=(5m-6)/3 === > 18m+9=35m-42 ===> 35m -18m = 9+42 ===> 17m = 51 ==> m=3